OpenStudy (anonymous):

write √ (3/4)^9 as a power

OpenStudy (jdoe0001):

\(\bf \sqrt{\left(\cfrac{3}{9}\right)^9}\\ \quad \\ \quad \\ \textit{keep in mind that } \large a^{\frac{n}{m}} = \sqrt[m]{a^n}\)